What is the integration of (x^2 + 20)/(x sinx +5cosx)^2?



Grade:

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
If you simply go by substitution method, you will get
I = \int \frac{x^{2}+20}{(xsinx + 5cosx)^{2}} = \frac{5sinx - xcosx}{xsinx+5cosx} + constant
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty